OpenStudy (anonymous):

During a rescue operation, a 53oo-kg helicopter hovers about a fixed point. The helicopter blades send air downward with a speed of 62 m/s. What mass of air must pass through the blades every second to produce enough thrust for the helicopter to hover?

OpenStudy (jamesj):

Remember that we can express change in momentum, the impulse as the integral of Force with respect to (wrt) time, \[ \Delta p = I = \int F \ dt \] That in turn means that we can also express Force as the derivative of p wrt time \[ F = \frac{dp}{dt} \] From this we can recover the usual expression, Newton's law, F=ma when mass is constant because \[ F = \frac{dp}{dt} = \frac{d \ }{dt} (mv) = m \frac{d \ }{dt} v = ma \] But in the case of the helicopter, what we are measuring is the impulse of the thrust from the air moved by the propeller, call it \( F_p \) and m is what is changing and v is constant, hence \[ F_p = \frac{d \ }{dt} (mv) = v \frac{dm}{dt} \] Now, if the aircraft is hovering, the force from the thrust of the propeller must exactly equal the gravitational force, \( F_g = Mg \), where M is the mass of the helicopter. Hence \( F_g = F_p \), i.e., \[ Mg = v \frac{dm}{dt} \] You now just need to solve for dm/dt.

OpenStudy (jamesj):

[ Or to put it in simpler terms without derivatives, Force x time = Impulse = Change in momentum = Change in (mass x velocity): \[ F \Delta t = I = \Delta p = v \Delta m \] Hence \[ F = v \frac{\Delta m}{\Delta t} \]